OpenStudy (anonymous):

Hey I was just wondering if someone could help with this question 2y+z=3-3x x-3y=4-z -6x-4y=2z+1 Directions are to solve the following system of equations and all answers must be in simplest form.

OpenStudy (amistre64):

where ya stuck on it?

OpenStudy (amistre64):

first thing to do is set it up so it looks normal to work with 3x +2y +z = 3 x -3y +z = 4 -6x -4y -2z = 1

OpenStudy (anonymous):

Well I first fixed the equations so variables would be on one side The i used the first and second equation to to try and elminate the variables but i ended up elminating all the variables somehow

OpenStudy (amistre64):

lets kill the xs lol

OpenStudy (anonymous):

ok lol

OpenStudy (amistre64):

or you wanna do the zs?

OpenStudy (amistre64):

zs look precarious, go in for the kill

OpenStudy (anonymous):

ok zz it is

OpenStudy (anonymous):

you would get -2x-5y=1 if you use the first and third equation

OpenStudy (amistre64):

3x +2y +z = 3 -x +3y -z = -4 -------------- 2x +5y = -1 -6x -4y -2z = 1 2x -6y +2z = 8 ---------------- -4x -10y = 9

OpenStudy (anonymous):

ok i get that then you do those two together right?

OpenStudy (amistre64):

2x +5y = -1 <- death to ys :) -4x -10y = 9 4x +10y = -2 -4x -10y = 9 ---------------- yeah, these lines are parallel to each other lol

OpenStudy (amistre64):

there is no solution to the system; it happens :)

OpenStudy (amistre64):

same slopes; different points...

OpenStudy (anonymous):

ok thank you soo much

OpenStudy (anonymous):

I thought there was something wrong when they all canceled

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

youre welcome :) and quit breaking the math...others have to use it lol

OpenStudy (anonymous):

ha ha

OpenStudy (anonymous):

I think i can figure the rest out now

OpenStudy (anonymous):

thanks again :)

OpenStudy (amistre64):

come back now ya hear....we got all the right answers :)

OpenStudy (anonymous):

def will it was a big help